OpenStudy (anonymous):

Given that f(x) = 5x − 10 and g(x) = x + 3, solve for f(g(x)) when x = −1. −30 0 10 12

OpenStudy (anonymous):

first, plug g(x) into f (x): ( f o g)(x)= f (g(x))=5(X+3)-10 so f (g(-1))=5(-1+3)-10=???

OpenStudy (anonymous):

can you do it?

OpenStudy (anonymous):

@Duck.

OpenStudy (anonymous):

Give me a minute.